29. In the small island nation of Carribeau, there are three economic agents. One is a company that does the fishing; second is a restaurant that makes sushi; and lastly there is the government. The fishing company collects 10 tons of fish and sells 8 tons to the restaurant and 2 tons to customers at $3 million per ton, and pays wages of $10 million and taxes of $5 million. The restaurant buys 8 tons of fish as noted before and has $50 million in revenue from selling sushi. The fish is the only cost and they pay $10 million for wages and $10 million for taxes. The government spends all its tax revenue on wages for its employees. Using the expenditure approach, what is the consumption of C+I+G+NX and what is the total GDP? a. C $80 million; GDP-$115 million b. C $56 million; GDP-$95 million c. C $80 million; GDP-$95 million d. C $56 million; GDP-$71 million30. Suppose employment in Blueland was 9,553,800 in December and 9,574,800 in January. What was the annualized monthly percentage change in employment between December and January? a. -021% b. 0.22% C. 122% d. 2.67%34. Last year the Republic of Econ had real GDP of 27.0 billion. This year it had real GDP of 31.5 billion. Which of the following changes in population is consistent with a5 percent growth rate of real GDP per person over the last year? a. The population decreased from 88 million to 84 million b. The population decreased from 75 million to 73 million c. The population increased from 45 million to 50 million d. The population increased from 60 million to 62 million
Answer #1

Q29) option D)

in expenditure method :

Consumption Spending on goods & services , C = 2*3 + 50

= 56

As 2 tons are sold to consumers at $3 million per ton

While 8 tons are sold to restaurant, ( which is not included)

& Restaurants finally sell for $ 50 million

Now GDP = C + G

As I & NX = 0

G = wages & salaries paid by govt

= 10+5 = 15

So GDP = 56+15 = 71

Q30) option D)

Annualised monthly % change=[ ( 9574800/9553800)12 - 1 ] * 100

= 2.67%

Q31) option C)

now growth rate of real GDP per capita = growth rate of real GDP - growth rate of population

Now growth rate of real GDP = [(31.5-27)/27]*100

= 16.66%

So population growth rate = 16.66-5

= 11.66 %

So in C) growth rate of population = (50-45)/45 = 11.11%